OpenStudy (anonymous):

@whpalmer4 4x^2+4x=15. solve by quadratic formula. does this mean I use the x=-b= +/-sq.rt b2-4ac/2a or do i use the ax^2 +bx +c =0 method?

OpenStudy (anonymous):

The first one\[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] is the quadratic formula

OpenStudy (anonymous):

thank you. give me a minute to pen and paper this to see if i can get a correct answer

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

okay the closest to an answer I can get is (-4 +/- 16)/8. not quite sure how to proceed to further simplify

OpenStudy (anonymous):

-20/8 or 12/8 then reduces to...-5/2 or 3/2, yes? so x= -5/2 or 3/2?

OpenStudy (anonymous):

Quite right!

OpenStudy (anonymous):

thankee-sai. your help is most appreciated :)

OpenStudy (anonymous):

No problem, you did all of the work :)

OpenStudy (anonymous):

i usually just need a little nudge to point me in the right direction. i won't learn a thing if someone else does the work

OpenStudy (anonymous):

I agree and its great that you realize that, especially with math. Feel free to tag me anytime you need a little nudge lol

OpenStudy (anonymous):

thankee-sai and will do. Blessed be!

OpenStudy (anonymous):

And you!

OpenStudy (whpalmer4):

You have to put it in the form \(ax^2+bx+c=0\) first in order to get the correct values of \(a,b,c\) for the quadratic formula \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (whpalmer4):

\[x=-\frac{5}{2},\,x =\frac{3}{2}\]are the correct solutions once you do that.